OpenStudy (anonymous):

Solve 3x – 5y = 17 9x – 15y = –4 (–1, –4) (4, –1) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

@ Hero

OpenStudy (anonymous):

use elimination i think

Parth (parthkohli):

Multiply the first equation by 3.\[9x - 5y = 51 \]

Parth (parthkohli):

Now, we have two equations.\[ 9x - 15y = 51\]\[ 9x -15y = -4 \]

Parth (parthkohli):

If we use substitution, then we got something with us.\[ 9x - 15y = 51 = -4\]Which implies\[51 = -4\]

Parth (parthkohli):

Now I don't know about other universes, but the solution does NOT apply on Earth.